OpenStudy (anonymous):

y''+y'+y = 2+x+cos x The question is find a particular integral for it, how do I begin?

OpenStudy (anonymous):

1/ find general solution 2/ find particular do 1 first.

OpenStudy (anonymous):

okey so the general solution is a*y'' +b*y' +c*y, a,b and c are constants and the sum of the particular integral and the complementary function is y = y(pi)+ y(cf) and y(cf) contains no arbitrary constants and y(ct) contains two arbitrary constants, how do go about?

OpenStudy (anonymous):

Characteristic equation for homogeneous part is r^2+r+1 =0 gives us roots r =\(-\frac{1}{2}\pm i\sqrt3\) --> the general solution is???

OpenStudy (anonymous):

I'm totally lost

OpenStudy (anonymous):

You have to do part 1 to find general solution because it tells you what particular solution is. That's why I ask you do that.

OpenStudy (anonymous):

okey so y = eâx(Acos Bx +Bsin Bx)?

OpenStudy (anonymous):

so that gives me e^-.5x(Acos 1,23x + B sin 2,23x)?

OpenStudy (anonymous):

In this case, the general solution has the form of \(y = e^{-x/2}(A cos (\sqrt3 x)+B sin(\sqrt 3 x)\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Now, non-homogenous part: what is the form of the solution??

OpenStudy (anonymous):

don't forget you have 2 parts: polynomial and trig for polynomial, do as usual for trig must times x to make it different from general solution

OpenStudy (anonymous):

okey Ill try to see if I can solve it now, hopyfully I'll get it right.

OpenStudy (anonymous):

Okey if y = e−x/2(Acos(3√x)+Bsin(3√x), and y = y(pi) + y(cf) What is y(pi) ?? @OOOPS

OpenStudy (anonymous):

it says y(pi) = 1+x+sinx, but how??

OpenStudy (anonymous):

I don't know your notations. But I guess y (pi) is y (general) and y(cf) is y (particular) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, for non-homogenous part, you have 2 parts : polynomial + trig 1/ for polynomial, let's call the solution of this part is y1 , find it out by solving y" +y'+y = 1+x the solution y1= ax^2+bx+c y1' = y"= -------------------------- you work on it

OpenStudy (anonymous):

Know how to do??

OpenStudy (anonymous):

okey so y1' is = 2ax+b but y''= that s the tricky part

OpenStudy (anonymous):

y"= 2a , why is it hard?

OpenStudy (anonymous):

did u mean y''=y1''?

OpenStudy (anonymous):

oh in that case it wasnt hard, I thought for something else okey, yeah but now im with you.

OpenStudy (anonymous):

no, hihi since y1 is solution, right? so that y1 satisfies the equation y1" + y1'+y1=1+x

OpenStudy (anonymous):

and from above, we replace all ax + 2ax+b+ax^2+bx+c= 1+x One more question: do you have initial values?? like y(0) = .... y(1) =...

OpenStudy (anonymous):

nope i dont.

OpenStudy (anonymous):

it just says find the particular integral.

OpenStudy (anonymous):

Wow!! so everything is on letter form?? I mean ax +b......

OpenStudy (anonymous):

yeah it goes like this Q = y''+y'+y = 2+x+cos x A = y(particular) = 1+ x+ sin x

OpenStudy (anonymous):

ok, back to the problem, got me so far from the ax+2ax +b+ax^2 +bx +c =1+x

OpenStudy (anonymous):

now how do we make that into 1 + x + sin x? (Im starting to feel so stupid)

OpenStudy (anonymous):

We MUST break it into 2 parts; I just solve the first part: polynomial

OpenStudy (anonymous):

from the above equation: we have a=0 3a+b=1 --> b =1 b+c=1 --> c =0 got it??

OpenStudy (anonymous):

ah yeah I think so, and where does the sin x come in?

OpenStudy (anonymous):

so, y1= ax^2+bx+c= x the first partial solution is x now the second part solve y2 = ecosx + fsinx y2'=? y2"= ? ---------------------------- ??

OpenStudy (anonymous):

okey. y2' = cos x but why fsinx tho? y2'' ? -sin x

OpenStudy (anonymous):

I use e, f as coefficients to differentiate from coefficients above. We used A, B, a, b,c already, I don't want to use them again. That's it

OpenStudy (anonymous):

straight up, this is so hard, because our test is in another language, so the terms and everything will be more confusing, we do everthihng in english and have exams in swedish...

OpenStudy (anonymous):

The standard form of partial solution for sin x, cos x, is Asinx + B cosx solve from it

OpenStudy (anonymous):

ah okey, then it goes -esin x + fcos x y'' is -ecos x -fsin x

OpenStudy (anonymous):

hihihi

OpenStudy (anonymous):

this may seem far fetched , I was an A student until I stopped caring...man its hard getting back to being so good..

OpenStudy (anonymous):

y2= esinx + fcosx y2'= ecosx-fsinx y2"= -esinx -fcosx ------------------------ y2"+y2'+y2= ecosx -fsinx= cosx which gives us e =1 f =0, --> so y2= cosx COMBINE y(partial) = y1+y2= x +cos x done

OpenStudy (anonymous):

hey, friend. I am vietnamese. When I study this stuff in America, I struggled a lot with my English and Math in English. hihihiihi

OpenStudy (anonymous):

thanks for the looong patience. maybe I'm tired or something, I'll take a break and hopefully get back on track. Thanks alot for all the help and patience :)

OpenStudy (anonymous):

ok.